OpenStudy (anonymous):

How would you solve 3onthesqrt(4x-4) +2 =0

OpenStudy (anonymous):

\[\sqrt[3]{4x-4} +2 = 0\]

OpenStudy (anonymous):

\[\sqrt[3]{4x-4} +2 = 0\] \[\sqrt[3]{4x-4} =-2\] \[(\sqrt[3]{4x-4})^3 =(-2)^3\] \[4x-4=-8\] \[4x=-4\] \[x=-1\]

OpenStudy (anonymous):

Thanks, but it wants to know how many solutions there would be as well, so would it be one or?

OpenStudy (anonymous):

only one solution

OpenStudy (anonymous):

thank you so much precalc homework.. it's about to be a long semester

OpenStudy (anonymous):

yw